OpenStudy (sambam):

Evaluate (x^-2y^0)/(x^3y^-2) when x = 2 and y = 5. a. 25/2 b. 25/32 c. 50

OpenStudy (anonymous):

\[\frac{y^2}{x^5}=\frac{5^2}{2^5}=\frac{25}{32} \]